### v1

# Language Modeling for the Future of Finance: A Quantitative Survey into Metrics, Tasks, and Data Opportunities

Abstract
Recent advances in language modeling have led to growing interest in applying Natural Language Processing (NLP) techniques to financial problems, enabling new approaches to analysis and decision-making. To systematically examine this trend, we review 374 NLP research papers published between 2017 and 2024 across 38 conferences and workshops, with a focused analysis of 221 papers that directly address finance-related tasks. We evaluate these papers across 11 qualitative and quantitative dimensions, identifying key trends such as the increasing use of general-purpose language models, steady progress in sentiment analysis and information extraction, and emerging efforts around explainability and privacy-preserving methods. We also discuss the use of evaluation metrics, highlighting the importance of domain-specific ones to complement standard machine learning metrics. Our findings emphasize the need for more accessible, adaptive datasets and highlight the significance of incorporating financial crisis periods to strengthen model robustness under real-world conditions. This survey provides a structured overview of NLP research applied to finance and offers practical insights for researchers and practitioners working at this intersection.
1 Introduction
Language modeling has become a key tool in natural language processing (NLP) for analyzing unstructured text, including reports, news articles, and social media posts. These methods are increasingly applied to finance-related tasks such as sentiment analysis (Balakrishnan et al., 2022), information extraction (Huang et al., 2023), summarization (Khanna et al., 2022), stock prediction (Jain & Agrawal, 2024), and volatility forecasting (Niu et al., 2023). Given the increasing number of NLP papers addressing financial problems, there is a need to systematically examine how the field is contributing to finance-related applications. While prior surveys have examined this intersection (Table 1), many adopt a qualitative approach (Chen et al., 2022b; Gao et al., 2021b; Xiao et al., 2024a), focusing on broad NLP techniques (Jagdale & Deshmukh, 2025; Man et al., 2019; Liu, 2024), deep learning (Ozbayoglu et al., 2020) and Large Language Models (LLMs) (Nie et al., 2024; Li et al., 2023c), or specific tasks such as sentiment analysis (Mishev et al., 2020). However, these studies often lack systematic evaluation.
In this study, we focus specifically on NLP research applied to finance. Our scope includes papers published in NLP venues where methods are developed, benchmarked, or adapted for financial problems. We analyze 374 papers published from 2017 to 2024 across 38 NLP conferences and workshops. After further filtering (Section 2), we retain 221 papers that directly address finance-related tasks. These are evaluated across 11 qualitative and quantitative dimensions, including tasks, methodologies, datasets, evaluation metrics, and accessibility.
Our analysis reveals several patterns in the current research landscape. Commonly addressed tasks include sentiment analysis, information extraction, and question answering, while areas such as explainability and privacy-preserving techniques are explored less frequently (Section 3). Standard machine learning metrics are widely used for evaluation, although these may not always reflect domain-specific priorities (Section 7.1). Most studies focus on stable market conditions, with less attention to more volatile periods (Section 7.2) and relying on outdated sources or overlooking issues such as survivorship bias (Section 7.3) Finally, we observe a trend toward the use of general-purpose language models over custom architecture (Section 4). While this reflects the rapid integration of recent advances in NLP, further exploration of domain-specific strategies may offer additional value for finance-related tasks.
Paper Review Years Papers Reviewed Features Area surveyed Systematic Collection Domain trends Quantitative Analysis Temporal Analysis
(Gao et al., 2021a) 1959 (A)-2020 87 5 General Overview × × × ×
(Liu, 2024) 2022-2024 49 11 General Overview × × × ×
(Millo et al., 2024) 2018-2023 30 1 Methodologies × × ×
(Chen et al., 2020b) 2016-2019 62 3 Financial Technology × × × ×
(Chen et al., 2022a) 2018-2022 38 2 Financial Technology × × ×
(Xing et al., 2017) 1998 (A)-2016 127 4 Financial Forecasting × × ×
(Zhao et al., 2024a) 2004 (A)-2024 146 1 Large Language Models × × ×
(Li et al., 2024) 2020-2023 68 1 Large Language Models × × × ×
(Dong et al., 2024) 2023-2024 206 2 Large Language Models × ×
(Nie et al., 2024) 2019-2024 318 1 Large Language Models × × ×
(Lee et al., 2025) 2018-2023 51 3 Large Language Models × × × ×
(Man et al., 2019) 2004-2019 89 1 Machine Learning × × × ×
(Ozbayoglu et al., 2020) 1998 (A)-2020 151 6 Deep Learning × ×
(Mishev et al., 2020) 2003 (A)-2020 89 1 Sentiment Analysis × ×
Language Modeling for the Future of Finance 2017-2024 374 11 Tasks, Methodologies, Data, Metrics, Code, Authorship, Funding
Table 1: Comparison of previous Natural Language Processing surveys in finance based on their focus areas, number of papers reviewed, and analytical features. Due to the absence of systematic collection methods in most prior research work, entries marked with (A) in the "Review Years" column indicate that the reported range is approximate.
2 Paper Extraction Process
Figure 1: Overview of our paper selection process and analysis dimensions. We collected papers from a broad range of NLP venues using abstract-level keyword filtering, yielding 374 candidates. After removing mismatches and shared task papers, we retained 221 papers, categorized into four groups based on their connection to financial tasks. The analysis focuses on the selected dimensions as shown in (Table 5, Appendix B).
To study how NLP has been applied to finance, we selected papers from 38 NLP venues, including CoLM, ACL, NAACL, EMNLP, LREC, COLING, and workshops like FinNLP and the Workshop on Economics and Natural Language Processing. The full list of venues is provided in Table 9 (Appendix E). As illustrated in Figure 1, we began by filtering papers that mentioned “finance” or “financial” in their abstracts. Out of all such papers published from 1975 to the present, more than 94% were published from 2017 onward. This growth coincides with the emergence of transformer-based models (Vaswani et al., 2017), which greatly expanded the scope of NLP applications. Accordingly, we use 2017 as a threshold year for our analysis.
Category Description
Financial Forecast (Category I) Targets predicting financial events, including stock movements, volatility, bankruptcy, and currency exchange rates.
Financial Resources (Category II) Covers tasks addressing finance-specific issues beyond prediction, such as constructing financial datasets, detecting fraud in finance-related documents, and extracting financial events.
Financial Applications (Category III) Focuses on general ML/NLP tasks like sentiment analysis and information/relation extraction for financial datasets.
Finance Related (Category IV) Covers tasks not applied to financial data or targeting financial problems, but potentially useful in finance, such privacy-preserving and explainable AI.
Table 2: Summary of task categories based on the specifics of their financial focus, ranging from direct prediction of financial outcomes to general NLP tasks with potential financial value.
This initial filtering returned 374 papers. Among them, 88 were tied to shared tasks (e.g., SemEval-2017 (Kar et al., 2017), FinCausal-2022 (Mondal et al., 2022)), and 65 used the term “financial” in a different contexts – such as referencing financial resources (Sekeres et al., 2024; Ding & Riloff, 2018) – without any actual financial application. After removing these, we retained 221 papers for our analysis.
Unlike previous surveys, which often emphasize qualitative observations, our study combines both qualitative and quantitative methods to provide a broader view of the field. We categorized the selected papers by their primary focus, resulting in four task-based categories (Figure 2). This classification helps us examine trends in how NLP techniques are being used to tackle finance-relevant problems and supports more detailed comparisons across types of tasks.
3 Task Distribution in NLP for Financial Applications
Figure 2: Distribution of primary tasks across categories. Each cell shows the task name and paper count (e.g., "Sentiment Analysis (3)"), with color gradients indicating frequency – darker shades represent more papers. "Miscellaneous" groups tasks that appear only once within Categories II and III.
To understand how NLP has been used in financial contexts, we categorized papers into four groups (shown in Table 2) based on their primary tasks (Figure 2). This classification helps map how language modeling techniques are applied and how directly each paper engages with finance, with Category I representing the most direct financial use cases. Detailed descriptions of each category are provided in Appendix D.
Financial Forecast (Category I) papers cover predictive tasks such as stock price and volatility forecasting. While these areas are well-studied, others like economic forecasting (Arno et al., 2023), risk assessment (Zhou et al., 2020), and cryptocurrency prediction (Seroyizhko et al., 2022) receive less focus. These areas present opportunities for expanding the reach of prediction models and improving their robustness and comparability (Section 7).
Financial Resources (Category II) papers often focus on dataset construction (Section 4, Figure 3). These datasets – ranging from annotated earnings calls to news and speeches – support tasks like financial event extraction (Huang et al., 2024; Ju et al., 2023), fraud detection (Erben & Waldis, 2024; Wang et al., 2019), and annotation (Aguda et al., 2024; Khatuya et al., 2024). Although these tasks contribute to enhancing financial data processing, other tasks remain relatively underrepresented.
In the Financial Applications (Category III) group, sentiment analysis (Rodriguez Inserte et al., 2023), information extraction (Lior et al., 2024), and question answering (Kosireddy et al., 2024) are the dominant tasks. These methods are widely used to analyze earnings calls, reports, and market commentary, where investor sentiment and factual extraction are key inputs for decision-making. Recent efforts have aimed to improve QA models for handling financial text (Theuma & Shareghi, 2024; Liu et al., 2024; Mavi et al., 2023).
Finally, among Finance-Related (Category IV) papers, the most studied areas are dataset construction and numerical reasoning. While these datasets are not finance-specific, they have potential applications in finance. For instance, fake news detection datasets (Vargas et al., 2021) can help reduce misinformation in markets. Numerical reasoning (Akhtar et al., 2023), important for understanding financial statements and assessing risk, is receiving more attention. However, other areas – such as explainable AI (XAI) (Klein & Walther, 2024) and privacy-preserving methods (Abbe et al., 2012) – remain rarely explored, despite their relevance to secure and interpretable decision-making (Basu et al., 2021).
4 Language Model Adoption & Datasets
Figure 3: Timeline of PLM/LLM adoption in NLP research applied to finance, alongside key datasets by task type: C (custom models), QA (question answering), IE (information extraction), OL (other language resources), and FD (financial documents). For each model, the top date indicates release; the bottom, first usage in the surveyed papers.
Acceleration of Model Adoption
Figure 3 (Language Models) shows how quickly language models are adopted for finance-related NLP tasks. While early models like BERT (Devlin et al., 2019) took 13 months to appear in these applications, recent models – GPT-3.5, GPT-4 (Achiam et al., 2023), and Mistral (Jiang et al., 2023) – were adopted within 6–12 months. LLaMA-2 (Touvron et al., 2023)) saw integration in just 5 months, reflecting the community’s increasing speed in applying general NLP advances to financial tasks.
Dominance of General-Purpose Models
As shown in Figure 3 (Language Models), general-purpose models – LLMs like GPT-4 and LLaMA-2, and Pretrained Language Models (PLMs) like BERT and RoBERTa (Liu et al., 2019)) – are used far more frequently than domain-adapted ones. BERT appears in over 70 papers, while finance-tuned models such as FinBERT (Araci, 2019; Yang et al., 2020b; Liu et al., 2020; Huang et al., 2022), FinMA (Xie et al., 2024a), and FinGPT (Yang et al., 2023) are less common. Notably, earlier FinBERT versions are more widely used than later ones, suggesting that improvements in general models often outweigh the benefits of domain-specific adaptation.
Decline of Custom Architectures
Custom models were common before 2022, especially for tasks like table-based QA (e.g., TagOp for TAT-QA (Zhu et al., 2021), HyBrider for HybridQA (Chen et al., 2020c)). Some models, like MDRM (Qin & Yang, 2019), HTML (Yang et al., 2020a), and HAN (Hu et al., 2021), are still used as baselines, particularly in volatility prediction (Niu et al., 2023; Mathur et al., 2022). However, the rise of general LLMs has led to a sharp decline in custom architectures. As a result of the accelerated development and deployment of these general purpose models, researchers increasingly rely on adapting pretrained models instead of building domain-specific ones from scratch.
Language Coverage in Datasets
The OL and FD sections of Figure 3 highlights the scarcity of non-English datasets for forecasting tasks such as return or volatility prediction. One exception is CMIN (Luo et al., 2023), which includes English and Chinese financial data. Most datasets use English-only sources like EDGAR filings (Maheshwari et al., 2022) and financial news (Xie et al., 2024b), which may limit model performance in non-English markets. Creating multilingual forecasting datasets would support broader applicability and reduce language bias, as also noted by Jørgensen et al. (2023).
Since 2017, research applying NLP to finance has shifted from custom models toward general-purpose pretrained architectures. New models are adopted faster than before, and general PLMs have largely replaced finance-specific ones. This preference also influences dataset use: most forecasting datasets are English-only, limiting generalization to global markets.
5 Methods Used in NLP for Finance
Figure 4: Trends in methodology adoption over time, showing stable use of traditional techniques and rapid growth in PLMs and LLMs. Less frequent methods are omitted for clarity.
To understand the methodological landscape, we grouped the approaches used in the surveyed papers into nine categories (Appendix C, Table 6). While the total number of papers has grown, methods like statistical NLP, conventional machine learning, and embeddings have remained mostly flat. In contrast, there has been a sharp increase in the use of deep learning, PLMs and LLMs, reflecting the field’s shift toward pretrained architectures.
Shift Toward PLMs and LLMs
As seen in Figure 4, the first three categories – statistical NLP, embeddings, and conventional ML – have persisted but have seen reduced usage as PLMs and LLMs take lead. The rise began with models like BERT in 2019, which significantly improved context handling in financial texts (Li et al., 2017; Chen et al., 2019; Ein-Dor et al., 2019). PLMs are now widely used for tasks such as summarization and entity recognition. Since 2023, LLMs like ChatGPT-3.5 have seen rapid adoption (Tang et al., 2023; Li et al., 2023a; Ni et al., 2023), especially for text generation (Guo et al., 2023b) and financial QA (Uluoglakci & Temizel, 2024). This transition has also brought more complex deep learning architectures used for multi-step reasoning and domain-specific challenges (Ma et al., 2024; Zhao et al., 2024b; Su et al., 2024).
Underexplored Techniques
While pretrained models dominate, several methods remain underused despite their promise. Statistical modeling, graph-based approaches, and knowledge graphs could improve model robustness and interpretability, particularly for applications like causality detection (Tabari et al., 2018) or market dynamics (Menzio et al., 2024). Emerging techniques like GraphRAG (Edge et al., 2024) highlight the potential of such methods, yet they appear in only a fraction of papers. Their limited presence suggests opportunities for expanding methodological diversity in future research.
6 Code & Data Accessibility
Figure 5: Temporal trends in code and dataset availability, highlighting the shift toward open-source practices and the growing accessibility of NLP resources for finance.
As the adoption of language models accelerates, the issue of open accessibility in NLP research, especially in finance, becomes more relevant. As shown in Figure 5, open-source practices have become more common in recent years. Before 2021, closed-source code dominated, but the trend has shifted toward sharing code and datasets. This move enhances transparency, reproducibility, and collaborative progress in NLP research applied to finance (Whited, 2023). Despite this, some papers still include inactive links or empty repositories, limiting reproducibility. At the same time, dataset availability has steadily increased since 2017, pointing to growing awareness around open data. Maintaining functional, up-to-date repositories remains essential to support meaningful benchmarking and model development.
7 Potential in Financial Forecasting
Figure 6: Distribution of evaluation metrics used in Category I papers. Most rely on ML-based metrics, while only a few financial metrics appear repeatedly.
Category I papers reveal several areas where forecasting models could be improved. Most studies use general ML metrics, leaving out finance-specific measures that better capture practical utility. Crisis periods are rarely considered, limiting insights into how models behave under stress. Finally, there is a strong U.S. and English-language bias in data, with limited adoption of global financial datasets. Expanding these areas of exploration could strengthen model robustness and support broader real-world applicability.
7.1 Potential in Financial Metrics
As shown in Figure 6, Category I papers mostly rely on ML metrics like accuracy, F1, or MSE (Sawhney et al., 2020b; Wu, 2020; Yangjia et al., 2022). While useful, these do not fully reflect financial performance. Only three financial metrics appear more than once – Sharpe Ratio, Maximum Drawdown, and Cumulative Return (defined in Appendix G).
Finance-specific metrics offer risk-adjusted insights that ML measures overlook. Sharpe Ratio, for example, evaluates return relative to risk (Zou et al., 2022). Incorporating such metrics would improve comparability across models and align research outcomes with practical needs (Zhang et al., 2024; Zou et al., 2022; Sawhney et al., 2021b).
7.2 Crisis Years for Stress Testing
Figure 7: Distribution of data years used in Category I papers, highlighting publication trends, recession periods, and key financial data milestones.
Figure 7 shows that data usage begins to rise after 1993, when electronic filings became publicly available, with another increase in 2005 following the HTML filing mandate (U.S. Securities and Exchange Commission, ). However, usage dips around 2009 and 2020 – years associated with the global financial crisis and COVID-19 pandemic. These periods are essential for evaluating model robustness, yet they remain underused, likely due to data availability challenges prior to 2015 (Bloomberg L.P., 2015).
Despite these constraints, crisis periods offer valuable testing grounds. Most models are trained on stable conditions, which may not reflect their behavior during volatility. Including data from crises enables stress testing (Investopedia Staff, 2025b) and helps build forecasting models that remain reliable under instability.
7.3 Data Sources
Commonly Used and Overlooked Data Sources
Most forecasting models use common financial sources: stock prices, SEC filings, financial news, earnings calls (e.g., MDRM (Qin & Yang, 2019), StockNet (Xu & Cohen, 2018)). But many valuable datasets are either rarely used (e.g., Federal Reserve reports (Menzio et al., 2024)) or completely absent from NLP research. Public resources like FRED (Federal Reserve Bank of St. Louis, 2024) and Fama-French (Fama & French, 2024) offer rich macroeconomic indicators that remain unexplored. Incorporating these datasets could help capture macroeconomic trends and market dynamics, improving model reliability.
Data Accessibility and Maintenance
Access limitations are also a concern. Legacy datasets like MDRM are outdated, and APIs such as Yahoo Finance’s have been discontinued. Others, like CRSP, require paid access. This limits reproducibility and restricts access to reliable, up-to-date data. There’s a clear need for open, regularly updated datasets applicable for financial forecasting.
8 Discussion & Conclusion
This paper presents a systematic, quantitative analysis of NLP research applied to finance. From an initial pool of 374 papers across major NLP venues (Appendix E), we curated a final set of 221 papers after filtering. We categorized them by primary task (Section 3) and assessed each across 11 qualitative and quantitative dimensions (Table 5). Table 3 summarizes key trends and identifies opportunities where language modeling research can further contribute to finance-related problems.
Criteria Trends and Observations Potential Opportunities and Recommendations
NLP Tasks in Finance Sentiment analysis, information extraction, and question answering are the most frequently addressed tasks. Forecasting is mostly centered around stock and volatility prediction. Underexplored areas such as explainability, privacy-preserving methods, and tasks like bankruptcy or cryptocurrency prediction present valuable directions for future work (Section 3).
PLM/LLM Adoption General-purpose PLMs and LLMs have largely replaced finance-specific and custom architectures. At the same time, the use of statistical NLP and conventional ML continues to decline (Section 4). Revisiting finance-specific models and exploring alternative methods such as graph-based learning or hybrid statistical models could offer improvements for financial tasks (Section 5).
Evaluation Metrics Most studies rely on ML metrics such as accuracy and MSE, which do not reflect financial performance. Incorporating financial metrics like Sharpe Ratio or Maximum Drawdown would improve the practical relevance and comparability of predictive models (Section 7.1).
Crisis Periods Few studies include data from financial crises such as 2008–2009 or 2020–2021, focusing instead on stable periods. Including data from volatile periods can support stress-testing and help build models more resilient to real-world fluctuations (Section 7.2).
Data Selection and Bias Most studies use English-language datasets and U.S.-based financial sources. While there is growing diversity in the types of data used (e.g., news, filings, social media), much of it is static, with limited updates or adaptation to changing financial contexts. Developing multilingual and globally representative datasets, along with maintaining and updating existing ones, would support better generalization and long-term applicability of models across financial domains (Section 7.3).
Open Accessibility Code and dataset sharing has become more common, especially after 2021, enhancing transparency and reproducibility. Open access should be a standard in NLP-for-finance research when legally feasible. Maintaining functional and up-to-date repositories is key for long-term reproducibility (Section 5).
Table 3: Summary of trends and research directions in NLP research applied to finance, highlighting areas of focus, methodological shifts, and key opportunities for future exploration.
Implications for NLP & LM Researchers
NLP research applied to finance has largely centered on a few core tasks – sentiment analysis, information extraction, and question answering – while underexploring others like explainability and privacy-preserving methods (Section 3). The rapid adoption of PLMs and LLMs has made it easier to apply language models, but has also led to reduced methodological variety (Section 4). Graph-based models, statistical techniques, and other alternatives remain underused despite their potential (Section 5). While open-source practices have become more common, researchers should be more careful about maintaining functional and up-to-date repositories to ensure long-term reproducibility and meaningful benchmarking (Section 5). In addition, adopting financial-domain metrics can make evaluations more meaningful and aligned with practical goals (Section 7.1). Including crisis periods in datasets (Section 7.2) and improving data selection practices (Section 7.3) can help ensure robustness and consistency. More diverse datasets, especially beyond English and U.S. markets, are also needed to broaden applicability (Section 7.3).
Considerations for Financial Practitioners
While NLP methods are increasingly used in financial contexts, several challenges remain. Many models are trained on static or incomplete datasets, leading to biases that may affect predictions (Section 7.3). Evaluations often omit real-world considerations such as market volatility and financial risk (Sections 7.2, 7.1). Incorporating financial benchmarks and testing models under stress conditions could improve their practical value. Continued progress in this direction can better align research outcomes with decision-making needs in financial settings.
Ethics Statement
This study does not assess the merit or quality of individual papers. We do not suggest that any category or method is inherently better than others. Our goal is to map the volume and distribution of research applying NLP methods to finance, and to offer a foundation for further study, without making judgments about the value of specific approaches.
Acknowledgments
We did not receive any specific funding for this work. We thank Arnav Hiray and Michael Galarnyk for their comments.
Appendix A Authorship and Funding
Based on the analysis we conducted, we also made a binary classification of the following categories:
  • Authors affiliated with academia
  • Authors affiliated with Industry
  • Funding from industry
  • Funding from government sources
  • Funding from academic sources
Figure 8: This figure displays the correlation between authors from industrial and academic backgrounds within various types of funding such as academic, industrial and governmental.
As seen in the figure 8, we see that there is no major correlation between funding and authorship. Thus it does not imply that if an author with either industrial or academic is present in the list of authors, there will necessarily be funding from the industry/academia/government.
Category Count
Papers with Academic and Industrial authors 78
Papers with only Academic authors 118
Papers with only Industrial authors 25
Papers with Industrial Funding 49
Papers with Governmental Funding 61
Papers with Academic Funding 35
Table 4: Summary of the count of authors and funding within each category.
Appendix B Various Dimensions surveyed
Exploration dimension Description
Primary task The main task addressed in the paper, such as volatility prediction (Wang et al., 2024).
Sub-Tasks Additional tasks that support the primary objective, like using sentiment analysis to enhance stock price predictions (Jain & Agrawal, 2024).
Methodology Techniques applied, ranging from traditional machine learning to deep learning and large language models. See Section 5.
Code and data availability Whether the paper provides open-source code or data, and whether those artifacts are accessible (e.g., active links).
Contribution What type of contribution the paper has such as a dataset, framework, model, evaluation, and so on.
Comparison/baselines The baselines the paper compares their work to.
Data source The types of data used, including financial reports, social media, news articles, or other financial documents.
Metrics Evaluation criteria, including standard ML metrics (accuracy, precision) and finance-specific metrics (Sharpe ratio).
Data years The time periods covered by the data, such as crisis years or more stable financial periods.
Authorship Indicates what affiliation the authors of the paper have (academic, industrial).
Funding source The origin of the research funding (academic, industrial, or governmental) if any.
Table 5: This table outlines the dimensions examined in our analysis of NLP research focused on the financial domain. Dimensions that led to remarkable findings are discussed in the main paper, while an analysis of authorship and funding is provided in Appendix A.
Appendix C Common Methods used
Methodology Description
Statistical NLP Methods like TF-IDF, Bag-of-Words, and n-grams focused on extracting statistical patterns from financial text.
Embeddings Techniques such as Word2Vec, GloVe, and custom embeddings designed to map financial terms into vector spaces, improving downstream tasks.
Conventional ML Algorithms like Logistic Regression, Support Vector Machines (SVM), and Decision Trees, often used for classification and risk prediction.
Deep Learning (DL) Custom neural network architectures tailored for specific financial tasks such as stock price forecasting.
PLMs Encoder-only models (e.g., BERT) and Encoder-Decoder models (e.g., T5) are used for tasks like document summarization and information extraction.
LLMs Models like GPT-3.5 and LLaMA-2 focus on text generation and understanding complex financial language.
Statistical Modeling Involves correlation analysis, Granger causality, etc., to understand relationships in financial data.
Graphs Uses graph structures to model interactions in financial systems.
Knowledge Graphs Integrating structured knowledge within financial tasks to enhance model performance and reliability.
Table 6: This table categorizes the key methodologies into nine groups.“PLMs” stands for Pretrained Language Models, while “LLMs” stands for Large Language Models.
Appendix D Task Description
The following Table 7 and Table 8 describe the task description with examples.
Primary Task Examples & Insights
Finance-Related (Categories IV)
Explainable AI (XAI) refers to methods that make machine learning outputs more interpretable and less of a "black box." These papers show how small input changes can drastically affect model explanations (Sinha et al., 2021) and highlight the instability of tools like LIME in sensitive applications (Burger et al., 2023). In finance, where decisions must be traceable and justifiable, clear and reliable explanation methods are essential for building trust in tasks like fraud detection and risk assessment.
Privacy-preserving methods protect sensitive data while still allowing model training and analysis. Examples include homomorphic encryption for text similarity (Kim et al., 2022), federated learning for distributed training without raw data exchange (Zhao et al., 2024d), and domain adaptation that shares only model parameters (Xiao et al., 2024b). These methods are especially important in finance, where institutions must work with private data while staying compliant and avoiding breaches.
Numerical reasoning involves working with numbers and structured data to solve problems or develop algorithms. This includes solving math word problems using declarative knowledge (Roy & Roth, 2018), learning numeracy through number embeddings (Duan et al., 2021), and pretraining models for verifying tabular claims (Akhtar et al., 2023). In finance, good numerical reasoning supports tasks like market prediction and risk evaluation by improving how models interpret and reason about numbers.
Financial Applications (Category III)
Sentiment analysis helps assess market mood and forecast financial movements using sources like news, social media, and reports. Studies analyze opinions in reports and news (Rodriguez Inserte et al., 2023), investor sentiment on social media (Guo et al., 2023a), and tone in financial texts (Choe et al., 2023). As markets become more influenced by public opinion, sentiment analysis plays a growing role in trading and analysis.
Information extraction (IE) and relation extraction focus on identifying entities and linking them to events or other entities. Examples include extracting financial events from documents (Zheng et al., 2019), signals in reports (Huang et al., 2023), and patterns in social media (Conforti et al., 2022). These tools help connect companies, individuals, and financial instruments to real-world developments (Liou et al., 2021), supporting real-time analysis and prediction.
Question Answering (QA) enables systems to find specific information in financial documents, reports, and databases. QA models support quick access to relevant facts (Mavi et al., 2023), and recent work focuses on calibrating large language models (LLMs) for finance-specific tasks (Zhao et al., 2024c; Theuma & Shareghi, 2024; Addlesee, 2024). This area emphasizes not only building QA tools but also improving their accuracy for use in high-stakes financial contexts.
Table 7: Overview of key primary tasks in Categories IV (Finance-Related) above the mid line and Category III (Financial Applications) below. Each entry summarizes the task’s role and relevance in NLP research applied to finance, with representative examples and practical insights.
Primary Task Examples & Insights
Financial Resources (Category II)
Dataset and Resource Construction involves building structured, labeled financial datasets for model development and evaluation. These papers focus on creating large-scale annotated resources (Chen et al., 2020a), including datasets like Tab-CQA (Liu et al., 2023) and ConvFinQA (Chen et al., 2022c) designed for reasoning over tables and multi-step numerical queries. Other works assemble corpora from Earnings calls (Pardawala et al., 2025), financial reports (Shah et al., 2022), company filings (Zmandar et al., 2022), news (Tang et al., 2023), and government documents (Shah et al., 2023), enabling downstream tasks like entity linking, numerical reasoning, or forecasting.
Fraud detection leverages domain-specific data and models to identify deceptive financial behavior across platforms. Erben & Waldis (2024) identifies financial scams on Instagram using a fine-tuned BERT model deployed via browser extension and REST API. Another approach detects identity fraud through interactive dialogue, employing knowledge graphs and reinforcement learning to expose inconsistencies in claimed personal data (Wang et al., 2019). These systems highlight how NLP architectures can protect users from financial manipulation.
Financial Forecast (Category I)
Stock Price and Volatility Prediction aims to forecast stock movements or market instability using text data. These models use historical data (Sawhney et al., 2021a), news (Ahbali et al., 2022), and event-driven signals (Sawhney et al., 2020a; Wu, 2020) to predict stock trends. Volatility prediction further explores market sensitivity by analyzing unstructured text like press releases or financial articles (Qin & Yang, 2019), helping traders anticipate fluctuations.
Risk Prediction involves identifying the likelihood and impact of adverse financial events, such as defaults or market disruptions. These models analyze unstructured data like earnings call transcripts (Sang & Bao, 2022), regulatory filings, and legal documents to detect early signals of financial risk (Li et al., 2023b). Applications include credit risk estimation, fraud detection, and systemic risk monitoring (Zhang et al., 2024).
Table 8: Descriptions of major primary tasks in Category II (Financial Resources) above the line and Category I (Financial Forecast) below. Each entry summarizes the task’s focus and contribution to NLP research applied to finance, with representative examples.
Appendix E Conference Proceedings
Category Conferences and Workshops
Financial Forecast (Category I) ACL, Australasian Language Technology Association Workshop, CCL, EACL, EcoNLP Workshop, EMNLP, EMNLP (Industry), Financial Technology and Natural Language Processing (FinNLP) Workshop, ICCL, LREC-COLING, NAACL, NAACL (Industry), SRW
Financial Resources (Category II) ACL, ACL (Industry), CoNLL, EACL, EcoNLP Workshop, EMNLP, EMNLP-IJCNLP, e-Commerce and NLP Workshop, Financial Narrative Processing and MultiLing Financial Summarisation Workshop, Financial Technology and Natural Language Processing (FinNLP) Workshop, GWC, ICON, ICCL, LREC, LREC-COLING, NAACL, NAACL (Industry), NLP4PI Workshop, Pattern-based Approaches to NLP in the Age of Deep Learning Workshop, SIGHUM Workshop
Financial Applications (Category III) ACL, ATALA, Bridging Human–Computer Interaction and Natural Language Processing Workshop, CCL, CoLM, Computational Approaches to Subjectivity, Sentiment and Social Media Analysis Workshop, DeeLIO Workshop, EACL, EcoNLP Workshop, EMNLP, EMNLP (Industry), EMNLP-IJCNLP, Financial Narrative Processing and MultiLing Financial Summarisation Workshop, Financial Technology and Natural Language Processing (FinNLP) Workshop, ICON, ICCL, INLG, LREC, LREC-COLING, NAACL, NAACL (Industry), Natural Legal Language Processing Workshop, News Media Content Analysis and Automated Report Generation Workshop, Pattern-based Approaches to NLP in the Age of Deep Learning Workshop, Safety4ConvAI Workshop, Structured Prediction for Natural Language Processing Workshop, TextGraphs
Finance Related (Category IV) ACL, ACL-IJCNLP, BlackboxNLP Workshop, EMNLP, Financial Technology and Natural Language Processing (FinNLP) Workshop, LREC-COLING, SRW, TrustNLP Workshop
Table 9: Conference and Workshop Categorization
Appendix F Papers per year
Year 2024 2023 2022 2021 2020 2019 2018 2017 2016 2015 2014 2012 2010 2008 2006 2002 2001 1976 1975
Number of Papers 127 53 90 32 57 15 14 21 6 1 5 1 3 3 1 1 1 1 1
Table 10: Number of papers satisfying our filtration criterion per year.
Appendix G Glossary
Metric Definition
Sharpe Ratio Measures how much return an investment gives for each unit of risk, compared to a risk-free asset (Investopedia Staff, 2025a).
Maximum Drawdown The biggest drop from a peak to a low point in a portfolio’s value before it recovers (Investopedia Staff, 2023b).
Cumulative Return The total profit or loss from an investment over time (Investopedia Staff, 2023a).
Table 11: Key Definitions of financial terms used in our paper.