### v2

# Language Modeling for the Future of Finance: A Survey into Metrics, Tasks, and Data Opportunities

Abstract
Recent advances in language modeling have led to growing interest in applying Natural Language Processing (NLP) techniques to financial problems, enabling new approaches to analysis and decision-making. To systematically examine this trend, we review 374 NLP research papers published between 2017 and 2024 across 38 conferences and workshops, with a focused analysis of 221 papers that directly address finance-related tasks. We evaluate these papers across 11 quantitative and qualitative dimensions, and our study identifies the following opportunities: (i) expanding the scope of forecasting tasks; (ii) enriching evaluation with financial metrics; (iii) leveraging multilingual and crisis-period datasets; and (iv) balancing PLMs with efficient or interpretable alternatives. We identify actionable directions for research and practice, supported by dataset and tool recommendations, with implications for both the academia and industry communities.
1 Introduction
Language modeling has become a key tool in natural language processing (NLP) for analyzing unstructured text, including reports, news articles, and social media posts. These methods are increasingly applied to finance-related tasks such as sentiment analysis (Balakrishnan et al., 2022), information extraction (Huang et al., 2023), summarization (Khanna et al., 2022), stock prediction (Jain and Agrawal, 2024), and volatility forecasting (Niu et al., 2023). Given the increasing number of NLP papers addressing financial problems, there is a need to systematically examine how the field is contributing to finance-related applications. While prior surveys have examined this intersection (Table 1), many lack a structured approach (Chen et al., 2022a; Gao et al., 2021b; Xiao et al., 2024), focusing on broad NLP techniques (Jagdale and Deshmukh, 2025; Man et al., 2019; Liu, 2024), deep learning (Ozbayoglu et al., 2020) and Large Language Models (LLMs) (Nie et al., 2024; Li et al., 2023), or specific tasks such as sentiment analysis (Mishev et al., 2020).
In this study, we focus specifically on NLP research applied to finance. Our scope includes papers published in NLP venues where methods are developed, benchmarked, or adapted for financial problems. We analyze 374 papers published from 2017 to 2024 across 38 NLP conferences and workshops. After further filtering (Section 2), we retain 221 papers that directly address finance-related tasks. These are evaluated across 11 quantitative and qualitative dimensions, including tasks, methodologies, datasets, evaluation metrics, and accessibility.
Table 1: Comparison of previous Natural Language Processing surveys in finance based on their focus areas, number of papers reviewed, and analytical features. Due to the absence of systematic collection methods in most prior research work, entries marked with (A) in the "Review Years" column indicate the approximate range.
Paper Review Years Papers Reviewed Features Area surveyed Systematic Collection Domain trends Quantitative Analysis Temporal Analysis
(Gao et al., 2021a) 1959 (A)-2020 87 5 General Overview × × × ×
(Liu, 2024) 2022-2024 49 11 General Overview × × × ×
(Millo et al., 2024) 2018-2023 30 1 Methodologies × × ×
(Chen et al., 2020a) 2016-2019 62 3 Financial Technology × × × ×
(Chen et al., 2022b) 2018-2022 38 2 Financial Technology × × ×
(Xing et al., 2017) 1998 (A)-2016 127 4 Financial Forecasting × × ×
(Zhao et al., 2024) 2004 (A)-2024 146 1 Large Language Models × × ×
(Li et al., 2024) 2020-2023 68 1 Large Language Models × × × ×
(Dong et al., 2024) 2023-2024 206 2 Large Language Models × ×
(Nie et al., 2024) 2019-2024 318 1 Large Language Models × × ×
(Lee et al., 2025) 2018-2023 51 3 Large Language Models × × × ×
(Man et al., 2019) 2004-2019 89 1 Machine Learning × × × ×
(Ozbayoglu et al., 2020) 1998 (A)-2020 151 6 Deep Learning × ×
(Mishev et al., 2020) 2003 (A)-2020 89 1 Sentiment Analysis × ×
Language Modeling for the Future of Finance 2017-2024 374 11 Tasks, Methodologies, Data, Metrics, Code, Authorship, Funding
Our analysis reveals not only valuable insights but also actionable directions for both research and practice with dataset and tool recommendations. Forecasting tasks, while well-established, leave room for exploration, especially in areas such as risk and macroeconomic prediction (Section 3). Financial evaluation metrics are gaining traction and could further improve the practical applicability of models (Section 4.1). The growing availability of temporally diverse (Sections 4.2, 4.3), multilingual and multimodal (Sections 4.4) datasets enables more robust, globally applicable models. Finally, while PLMs and LLMs dominate, lightweight and interpretable methods remain viable for practical deployment (Section 5).
2 Paper Extraction Process
To study how NLP has been applied to finance, we selected papers from 38 NLP venues, including ACL, NAACL, EMNLP, LREC, CoLM, COLING, and workshops like FinNLP and the Workshop on Economics and Natural Language Processing. As illustrated in Figure 1, we began by filtering papers that mentioned “finance” or “financial” in their abstracts. Out of all such papers published from 1975 to the present, more than 94% were published from 2017 onward. This growth coincides with the emergence of transformer-based models (Vaswani et al., 2017), which greatly expanded the scope of NLP applications. Accordingly, we use 2017 as a threshold year for our analysis. The keyword filtering approach enabled us to cast a wide net, allowing for a high volume of papers. We were then able to filter these further using manual methods.
This initial filtering returned 374 papers. Among them, 88 were tied to shared tasks (e.g., SemEval-2017 (Kar et al., 2017), FinCausal-2022 (Mondal et al., 2022)), and 65 used the term “financial” in a different contexts (such as referencing financial resources (Sekeres et al., 2024; Ding and Riloff, 2018)) without any actual financial application. After removing these, we retained 221 papers for our analysis.
Unlike previous surveys, which often emphasize qualitative observations, our study combines quantitative and qualitative methods for a broader view of the field. We categorized the selected papers by their primary focus into four task-based categories (Figure 2). This classification helps us examine trends in how NLP techniques are used to tackle finance-relevant problems and supports more detailed comparisons across types of tasks.
3 Task Distribution in NLP for Financial Applications
To understand how NLP has been used in financial contexts, we categorized papers into four distinct groups (shown in Table 2) based on their primary tasks (Figure 2). This classification provides a mapping from language modeling techniques to their applications and the level of engagement with finance, with Category I papers representing direct financial use cases.
Table 2: Summary of task categories based on the specifics of their financial focus, ranging from direct prediction of financial outcomes to general NLP tasks with potential financial value.
Category Description
Financial Forecast (Category I) Targets predicting financial events, including stock movements, volatility, bankruptcy, and currency exchange rates.
Financial Resources (Category II) Covers tasks addressing finance-specific issues beyond prediction, such as constructing financial datasets, detecting fraud in finance-related documents, and extracting financial events.
Financial Applications (Category III) Focuses on general ML/NLP tasks like sentiment analysis and information/relation extraction for financial datasets.
Finance Related (Category IV) Covers tasks not applied to financial data or targeting financial problems, but potentially useful in finance, such privacy-preserving and explainable AI.
Figure 1: Overview of our paper selection process and analysis dimensions. We collected papers from a broad range of NLP venues using abstract-level keyword filtering, yielding 374 candidates. After removing mismatches and shared task papers, we retained 221 papers, categorized into four groups by their connection to financial tasks.
Financial Forecast (Category I) papers cover predictive tasks such as stock price and volatility forecasting. While these tasks are well-studied, areas like economic forecasting (Arno et al., 2023), risk assessment (Zhou et al., 2020), and cryptocurrency prediction (Seroyizhko et al., 2022) receive less focus. These areas offer opportunities to expand the reach of predictive models, improving robustness and comparability (Section 4).
Financial Resources (Category II) papers often focus on dataset construction (Figure 6). Ranging from annotated earnings calls to news and speeches, These datasets support tasks like financial event extraction (Huang et al., 2024; Ju et al., 2023), fraud detection (Erben and Waldis, 2024; Wang et al., 2019), and annotation (Aguda et al., 2024; Khatuya et al., 2024). While these tasks enhance financial data processing, other tasks remain relatively underrepresented.
In the Financial Applications (Category III) group, sentiment analysis (Rodriguez Inserte et al., 2023), information extraction (Lior et al., 2024), and question answering (Kosireddy et al., 2024) are the dominant tasks. These methods are widely used to analyze earnings calls, reports, and market commentary, where investor sentiment and factual extraction are key inputs for decision-making. Recent efforts have aimed to improve QA models for handling financial text (Theuma and Shareghi, 2024; Liu et al., 2024; Mavi et al., 2023).
Finally, among Finance-Related (Category IV) papers, the most studied areas are dataset construction and numerical reasoning. While these datasets are not finance-specific, they have potential applications in finance. For instance, fake news detection datasets (Vargas et al., 2021) can help reduce misinformation in markets. Numerical reasoning (Akhtar et al., 2023), important for understanding financial statements and assessing risk, is receiving more attention. However, other areas – such as explainable AI (XAI) (Klein and Walther, 2024) and privacy-preserving methods (Abbe et al., 2012) – remain rarely explored, despite their relevance to secure and interpretable decision-making (Basu et al., 2021).
Figure 2: Distribution of primary tasks across categories. Each cell shows the task name and paper count (e.g., "Sentiment Analysis (3)"), with color gradients indicating frequency – darker shades represent more papers. "Miscellaneous" groups tasks that appear only once within Categories II and III.
4 Potential in Financial Forecasting
Category I papers reveal several areas where forecasting models could be improved. Most studies use general ML metrics, leaving out finance-specific measures that better capture practical utility (Section  4.1). Crisis periods are rarely considered, limiting insights into how models behave under stress (Section  4.2). Finally, there is a strong U.S. and English-language bias in data, with limited adoption of global financial datasets (Section  4.3,  4.4). Expanding these areas of exploration could strengthen model robustness and support broader real-world applicability.
4.1 Advancing Evaluation with Finance Specific Metrics
As shown in Figure 3, Category I papers mostly rely on ML metrics such as accuracy, F1, or MSE (Sawhney et al., 2020; Wu, 2020; Yangjia et al., 2022). While useful, these do not fully reflect financial performance. Finance-specific metrics offer risk-adjusted insights that ML measures overlook. Incorporating such metrics would improve comparability across models and align research outcomes with practical needs (Zhang et al., 2024; Zou et al., 2022; Sawhney et al., 2021).
Consider two models that predict daily stock returns. Model A achieves a lower mean squared error (MSE) than Model B, yet when deployed in a trading strategy, it results in frequent trades, high volatility, and poor performance during market downturns. In contrast, Model B, despite its higher MSE, generates a smoother equity curve, experiences lower drawdown, and delivers a higher Sharpe Ratio thus making it more appealing from an application perspective. This highlights a critical insight emphasized in financial machine learning literature: traditional statistical metrics like MSE are often inadequate for evaluating trading strategies, as they fail to capture real-world profitability and risk exposure (Bailey et al., 2015; de Prado, 2018). Instead, financial metrics such as Sharpe Ratio, Maximum Drawdown, and Cumulative Return are essential because they reflect the model’s risk-adjusted performance and robustness across market regimes (Lo, 2002; Krauss et al., 2017; Takahashi et al., 2009) as well as the applicability of the model to practical real world situations.
Key Financial Metrics
There are certain financial metrics that measure models’ performance in a financial setting and provide insight into the their applicability. Sharpe Ratio quantifies risk-adjusted return by comparing excess returns to volatility (Investopedia, 2025a). Maximum Drawdown (MDD) captures the largest loss from peak to trough, indicating downside risk and model robustness (Hayes, 2024). Cumulative Return measures total profit over a period: a direct benchmark for model outputs (Investopedia, 2007).
Tools for Metric Calculation
Several Python libraries support quantitative finance and model-driven investment strategies. PyPortfolioOpt (Martin, 2021) offers portfolio optimization tools including Sharpe Ratio and risk-return metrics. QuantLib (QuantLib Team, 2000) provides analytics for pricing, drawdown, and custom risk modeling. Backtrader (Backtrader Team, 2015) enables backtesting of trading strategies with performance evaluation, ideal for simulating NLP-driven trading. bt (Morissette, 2025) supports rapid prototyping and comparison of forecasting models with built-in financial metrics.
4.2 Strengthening Robustness by Incorporating Crisis Periods
Figure 4 shows that researchers use financial data dating back to 1993, when electronic filings became publicly available, with another rise in 2005 after the HTML filing mandate (U.S. SEC, 2025). However, most publications use post-2009 (the global financial crisis) and pre-2020 (the COVID-19 pandemic) data, thus missing the crisis periods in evaluation. Crisis periods are essential for evaluating model robustness, allowing stress testing (Investopedia, 2025b) and helping build forecasting models that remain reliable under instability.
Consider the case of Long-Term Capital Management (LTCM). Its quantitative models were highly leveraged and calibrated almost exclusively on stable market data, failing to account for extreme tail events. When the 1997 Asian and 1998 Russian crises struck, LTCM’s Value-at-Risk (VaR) models severely underestimated exposure, leading to large-scale losses as spreads exploded and liquidity evaporated. Within weeks, LTCM lost over 90% of its equity, triggering systemic risk concerns across the global financial system (Department of Land Economy, 2003; Federal Reserve History, 2025; Vanity Fair Staff, 1998). As noted in academic and regulatory analyses, this collapse illustrates the dangers of relying solely on backtests from stable periods without incorporating the possibility of rare but catastrophic shocks.
Most current models are trained in stable conditions that may not reflect their behavior during volatility. While there were data availability challenges prior to 2015 (Bloomberg L.P., 2015), the crisis years remain underused. In Section 4.3, we highlight valuable data sources, including those with crisis years data.
4.3 Opportunities in Data Coverage
Most forecasting models use common financial sources: stock prices, SEC filings, financial news, earnings calls. But many valuable datasets are either rarely used (e.g., Federal Reserve reports (Shah et al., 2023; Menzio et al., 2024)) or completely absent from NLP research (e.g., shareholder letters). Public resources like FRED (Federal Reserve Bank of St. Louis, 2024) and Fama-French (Fama and French, 2024) offer rich but underexplored macro indicators. Incorporating these datasets could help capture macroeconomic trends and market dynamics, improving model reliability. Below, we highlight sources of valuable financial data, including those already used by the community and the other beneficial ones.
Figure 3: Distribution of evaluation metrics used in Category I papers. Most rely on ML-based metrics, while only a few financial metrics appear repeatedly.
Regulatory Filings and Earnings Calls
U.S. public companies file a range of documents such as 10-K, 10-Q, 8-K, credit agreements, proxy statements, S-1, S-3, and others are accessible via the SEC’s EDGAR platform, through free API calls and RSS feeds (U.S. SEC, 2024). While similar registries exist (e.g., Canada’s SEDAR, Japan’s EDINET, UK’s Companies House), they are rarely used, reinforcing a U.S.-centric and English-language bias (Section 4.4). Earnings calls, where firms discuss quarterly performance with investors, are another rich data source. Seeking Alpha transcribes  4,500 calls quarterly (Seeking Alpha, 2025); other providers include The Motley Fool (The Motley Fool, 2025), FactSet (FactSet Insight, 2025), Thomson Reuters (Thomson Reuters, 2025), and S&P Global (S&P Global Market Intelligence, 2025), though often behind paywalls. Companies also post transcripts/slides on IR websites. Public datasets include transcriptions and audio, such as S&P 500 calls from 2017 (Qin and Yang, 2019) and 1,213 companies’ data covering 2015-2018 (Li et al., 2020a).
Figure 4: Data year distribution in financial forecasting papers, annotated with major financial events and infrastructure milestones. Highlights underuse of crisis periods despite their importance for model robustness.
Financial News and Social Media
Real-time news from sources like Reuters, Bloomberg, Dow Jones, CNBC, FT, and WSJ is vital for forecasting. The Reuters-21578 dataset (UC Irvine, 2023), with 10,369 articles from 1987, remains a benchmark in text analysis. Larger archives such as FNSPID (Zhou et al., 2024) link 15.7 million articles (1999-2023) to S&P 500 stocks. News is accessible via APIs (e.g., Bloomberg (Bloomberg L.P., 2025a)) and free sources like Yahoo Finance (Yahoo Finance, 2025) and Google News (Google News, 2025). On social media, “FinTwit” and platforms like X and StockTwits are used for sentiment analysis. Key datasets include Sentiment140 (Kaggle Contributor, 2025; Chaudhary, 2020), StockNet (Xu and Cohen, 2018), and Reddit datasets (Kaggle Contributor, 2021; Wang et al., 2024). X’s API offers limited free and full historical access under paid tiers (X (2025), formerly Twitter). StockTwits, with Bullish/Bearish sentiment tagging, offers free and commercial API access (StockTwits, 2025) and is often used in forecasting studies.
Macroeconomic and Monetary Data
Central bank communications (policy statements, minutes, and press releases) offer critical macroeconomic guidance. The U.S. Federal Reserve’s FOMC releases detailed statements and transcripts on interest rates, inflation, and economic outlook (Federal Reserve System, 2025), with archives available on the Fed’s website. Despite their market relevance, such texts remain underused in NLP-based forecasting. Shah et al. (2023) addresses this by introducing an annotated dataset of FOMC communications and proposing a sentence-level classification task (hawkish, dovish, neutral) using fine-tuned transformers. Economic indicators like GDP, CPI, and unemployment figures – issued by the BEA (U.S. BEA, 2025), FRED (Federal Reserve Bank of St. Louis, 2024), Eurostat, IMF, OECD, and the World Bank (The World Bank, 2025) – are mostly numerical but often include press commentary, all freely available via official portals.
Analyst Reports
These are proprietary equity research reports with buy/sell calls and sector outlooks. Data providers such as Thomson Reuters, FactSet, Bloomberg, and S&P I/B/E/S aggregate these reports, typically through subscription services. All-in-one platforms like the Bloomberg Terminal (Bloomberg L.P., 2025b) and S&P Capital IQ (S&P Global, 2025) provide access to equity research, earnings estimates, and corporate event data as part of their premium offerings. There is no comprehensive free source; however, excerpts or key takeaways occasionally surface in news articles or on platforms like Seeking Alpha.
4.4 Data Concerns and Biases
Data Accessibility Concerns
Despite the presence of a large variety of data sources, there are significant limitations in data accessibility. Legacy datasets described in Section 4.3 often become outdated, and APIs for accessing data may be discontinued, as the Yahoo Finance API (EODHD, 2024), or restricted by paywalls. Other essential databases, such as CRSP, require paid access and are unavailable to most individual researchers. These barriers severely limit reproducibility and restrict access to reliable, up-to-date data. As a result, there is a clear need for open, regularly updated financial datasets suitable for forecasting.
Data Language and Country Bias
As shown in the previous sections, financial NLP research is heavily skewed toward English-language data, particularly from U.S. markets. Most forecasting datasets rely on U.S. regulatory filings, earnings calls, and financial news, which is often centered on large-cap firms like those in the S&P 500. This narrow focus limits the applicability of models to non-English and non-U.S. contexts. Figure 6 highlights that while some multilingual datasets exist, they are mostly designed for language modeling rather than financial forecasting. Very few datasets enable forecasting tasks in other languages, with rare exceptions like the bilingual CMIN corpus (Luo et al., 2023). Expanding multilingual, market-diverse datasets is essential for building globally robust forecasting models and mitigating systemic biases in NLP for finance, as also noted by (Jørgensen et al., 2023).
Multimodal Data Scarcity
Despite significant progress in multimodal learning, the domain of NLP in Finance still lacks comprehensive datasets that integrate text, audio, and other critical financial modalities. One of the few existing resources is MAEC, a dataset aligning transcripts with audio from over 3,400 S&P 500 earnings calls, totaling more than 920 hours of speech, specifically designed for financial risk forecasting (Li et al., 2020b). More recently, FinAudio has emerged as the first standardized benchmark focusing on audio-first tasks (ASR on short and long earnings call segments, plus summarization), compiling over 400 hours of financial audio from multiple sources to evaluate Audio-LLM capabilities (Cao et al., 2025). While many existing datasets support exploration of acoustic and textual cues in financial forecasting, truly multimodal resources (those integrating visual data such as stock charts, presentation slides alongside structured financial metrics) remain very limited. Notable exceptions include Mathur et al. (2022b) and Fons et al. (2024), which begin to address this gap. This modality gap highlights the pressing need for richer, tri- or quad-modal datasets to better capture the multifaceted signals that influence market dynamics.
5 Broader Insights for NLP in Finance
5.1 Complementing Foundation Models with Practical Alternatives
Based on our observations from our corpus of research papers, the methodological landscape in NLP for Finance has shifted decisively toward pretrained language models (PLMs) and large language models (LLMs). We group existing approaches into several broad categories, which allows us to track adoption trends and the rise of PLMs, beginning with the use of BERT-based models in finance around 2019 (Araci, 2019a; Yang et al., 2020c), and accelerating rapidly with the emergence of general-purpose LLMs like ChatGPT (Wu et al., 2023; Gao et al., 2023). However, NLP tasks in Finance often pose domain-specific constraints such as interpretability, regulatory transparency, data scarcity, and the need for low-latency systems. These factors make non-PLM methods (such as rule-based classifiers, feature-engineered models, and domain-specific embeddings) not only viable but often preferable in production environments (Shah et al., 2023; Lopez et al., 2021). The rest of this section revisits key methodological alternatives to PLMs, each addressing practical and conceptual needs that remain critical in industrial financial tasks.
Latency and Scalability
Classical NLP pipelines are far more resource-efficient than LLMs. In time-time contexts like news analytics or alert systems, lightweight models (e.g., logistic regression or small CNNs) deliver insights instantly and can scale to massive corpora without requiring GPU infrastructure (Zhai et al., 2019). Systems like RavenPack (Cornez, 2018) and Refinitiv News Analytics (TRNA) (Waeosri, 2020) use rule-based NLP and horizontally scalable architectures to deliver structured news data with sub-second latency, supporting real-time trading and high-throughput analysis.
Interpretability and Compliance
Lexicon- and rule-based methods remain essential in regulated settings. Financial sentiment lexicons (e.g., Loughran-McDonald) (Loughran and McDonald, 2011) and keyword rules are transparent, auditable, and often outperform neural models in compliance or risk-flagging tasks (Hosseini et al., 2018). Lexicon-based surveillance tools, such as those used in communications monitoring (NICE Actimize, 2023) and financial crime compliance (Risk & Compliance, 2021), are widely adopted for their transparency and ease of audit, allowing firms to justify alerts with specific keywords.
Information Extraction and Knowledge Graphs
Named-entity and relation extraction systems power knowledge graphs and schema-driven applications. These systems offer precision, structure, and explainability, which are ideal for fraud detection, Know Your Customer (KYC), and regulatory alignment, where PLMs often lack controllability (Szarvas et al., 2007; Tarnopolski et al., 2019). Financial institutions use rule-based NLP for entity and relation extraction to build and update knowledge graphs in KYC and Anti-Money Laundering (AML) workflows, as seen in solutions from KYCHub (2025) and TigerGraph (2024), enabling automated risk monitoring and explainable link discovery.
Data Efficiency and Domain Adaptability
Feature-based pipelines using TF-IDF, static embeddings, and classical classifiers (SVMs, XGBoost) work well with limited data. They enable explicit domain knowledge integration (such as tagging financial terms or numerical cues) for credit scoring, bankruptcy prediction, and event detection (Li et al., 2018; Khandani et al., 2021; Alanis et al., 2022). Financial firms often use TF-IDF (Simha, 2021) and XGBoost (Kamau, 2024) to build effective models with limited labeled data without requiring costly infrastructure.
Structure and Summarization
Unsupervised techniques like LDA and extractive summarization remain valuable for discovering latent topics or summarizing dense financial documents. These methods are fast, interpretable, and competitive on factual corpora like earnings calls or filings (Blei et al., 2003; Nenkova and McKeown, 2011). Companies like BlackRock, Bank of America (Phoenix Strategy Group, 2025), and American Express (Arora and Radhakrishnan, 2020) apply topic modeling and extractive summarization to earnings calls and filings, using unsupervised or minimally supervised methods for reliable and traceable outputs.
5.2 Implications of Shift Toward General-Purpose Models
General-purpose language models have rapidly become the default choice in NLP applied to Finance, with models like RoBERTa (Liu et al., 2019), GPT-4 (Achiam et al., 2023), and LLaMA-2 (Touvron et al., 2023) adopted far more widely and quickly than domain-specific alternatives. As Figure 6 shows, even earlier versions of FinBERT (Araci, 2019b; Yang et al., 2020b; Liu et al., 2020; Huang et al., 2022) are used more than newer ones, and recent finance-tuned models like FinGPT (Yang et al., 2023) remain rare. This trend reflects how easy it has become to apply strong general models out of the box. For practitioners, this means many models are optimized for benchmarks, not for the realities of finance, where interpretability, latency, or regulation often matter more than raw accuracy.
Figure 5: Trends in code and dataset availability, highlighting the shift toward open-source practices and the growing accessibility of NLP resources for finance.
Figure 6: Timeline of PLM/LLM adoption in NLP research applied to finance, alongside key datasets by task type: C (custom models), QA (question answering), IE (information extraction), OL (other language), and FD (financial documents). For each model, the top date indicates release; the bottom, first usage in the surveyed papers.
Table 3: Summary of trends and research directions in NLP research applied to finance, highlighting areas of focus, methodological shifts, and key opportunities for future exploration.
Criteria Trends and Observations Potential Opportunities and Recommendations
NLP Tasks in Finance Sentiment analysis, information extraction, and question answering are the most frequently addressed tasks. Forecasting is mostly centered around stock and volatility prediction. Underexplored areas such as explainability, privacy-preserving methods, and tasks like bankruptcy or cryptocurrency prediction present valuable directions for future work (Section 3).
Evaluation Metrics Most studies rely on ML metrics such as accuracy and MSE, which do not reflect financial performance. Incorporating financial metrics like Sharpe Ratio or Maximum Drawdown would improve the practical relevance and comparability of predictive models (Section 4.1).
Crisis Periods Few studies include data from financial crises such as 2008–2009 or 2020–2021, focusing instead on stable periods. Including data from volatile periods can support stress-testing and help build models more resilient to real-world fluctuations (Section 4.2).
Data Selection and Bias Most studies use English-language datasets and U.S. financial sources. While there is growing diversity in the types of data used (e.g., news, filings, social media), much of it is static, with limited updates or adaptation to changing financial contexts. Developing multilingual and globally representative datasets (Section 4.4), along with maintaining and updating existing ones (Section 4.3), would support better generalization and long-term applicability of models across financial domains.
PLM/LLM Adoption General-purpose PLMs and LLMs have largely replaced finance-specific and custom architectures. At the same time, the use of statistical NLP and conventional ML continues to decline. Revisiting finance-specific models and exploring alternative methods such as graph-based learning or hybrid statistical models could offer improvements for financial tasks (Sections 5.1, 5.2).
Open Accessibility Code and dataset sharing has become more common, especially after 2021, enhancing transparency and reproducibility. Open access should be a standard in NLP-for-finance research when legally feasible. Maintaining functional and up-to-date repositories is key for long-term reproducibility (Section 5.3).
This shift is also mirrored in the decline of custom architectures. Prior to 2022, task-specific models were more common: TagOp for TAT-QA (Zhu et al., 2021), HyBrider for HybridQA (Chen et al., 2020b)); MDRM (Qin and Yang, 2019), HTML (Yang et al., 2020a), and HAN (Hu et al., 2021) in volatility prediction (Niu et al., 2023; Mathur et al., 2022a)). Today, these efforts are rarely pursued, which seems like an efficiency gain at first glance. However, as Section 5.1 discusses, real-world financial applications often require design trade-offs, such as transparency for compliance, or fast inference for deployment, and off-the-shelf LLMs do not accommodate them well. The move away from custom models narrows the solution space at a time when financial tasks still demand methodological flexibility.
5.3 Enabling Reproducibility through Open Resources
As the adoption of language models accelerates, the issue of open accessibility in NLP research, especially in finance, becomes more relevant. As shown in Figure 5, open-source practices have become more common in recent years. Before 2021, closed-source code dominated, but the trend has shifted toward sharing code and datasets.
This move enhances transparency, reproducibility, and collaborative progress in NLP research applied to finance (Whited, 2023). Despite this, some papers still include inactive links or empty repositories, limiting reproducibility. At the same time, dataset availability has steadily increased since 2017, pointing to growing awareness around open data. Maintaining functional, up-to-date repositories remains essential to support meaningful benchmarking and model development.
6 Conclusion & Discussion
Drawing on Table 3, our survey shows that sentiment analysis, information extraction, and forecasting have driven NLP in finance, yet critical gaps remain: finance-specific evaluation metrics, stress-testing on crisis data, and truly global, multilingual datasets. These findings underscore both the field’s progress and the work ahead to build resilient, transparent, and inclusive financial NLP solutions. Below, we discuss their implications for researchers and practitioners.
Implications for NLP and Language Modeling Researchers
The application of LMs to financial domains presents challenges that differ from traditional NLP benchmarks. Our findings suggest that domain-specific requirements, such as interpretability, latency, and compliance, make classical or hybrid approaches highly relevant, even amid the LLM wave. The lack of financial evaluation metrics, underuse of crisis-period data, and limited attention to multilingual and global financial corpora signal areas where NLP researchers can contribute meaningfully. Moreover, the scarcity of domain-adapted models points to missed opportunities for more effective domain transfer. Future work should prioritize robustness, transparency, and real-world relevance over benchmark performance alone.
Considerations for Financial Practitioners
For finance professionals and institutions considering NLP solutions, this survey provides guidance on selecting appropriate tools. Off-the-shelf LLMs may appear attractive, but their performance in finance-specific contexts is often limited by lack of customization, interpretability, or latency guarantees. Classical and lexicon-based methods, when tailored to regulatory or operational constraints, can outperform black-box models in compliance or auditing settings. Practitioners should also evaluate model performance using finance-aligned metrics and test under historical stress scenarios. Finally, leveraging multilingual datasets and broadening market coverage can help mitigate geographic and systemic biases in automated decision-making.